107.     Porter Company purchased equipment for $450,000 on January 1, 2007, and will use the double-declining-balance method of depreciation. It is estimated that the equipment will have a 3-year life and a $20,000 salvage value at the end of its useful life. The amount of depreciation expense recognized in the year 2009 will be

a.   $50,000.

b.   $30,000.

c.   $54,440.

d.   $34,440.

108.     A plant asset was purchased on January 1 for $50,000 with an estimated salvage value of $10,000 at the end of its useful life. The current year's Depreciation Expense is $5,000 calculated on the straight-line basis and the balance of the Accumulated Depreciation account at the end of the year is $25,000. The remaining useful life of the plant asset is

a.   10 years.

b.   8 years.

c.   5 years.

d.   3 years.

109.     Equipment was purchased for $60,000. Freight charges amounted to $2,800 and there was a cost of $8,000 for building a foundation and installing the equipment. It is estimated that the equipment will have a $12,000 salvage value at the end of its 5-year useful life. Depreciation expense each year using the straight-line method will be

a.   $14,160.

b.   $11,760.

c.   $9,840.

d.   $9,600.

110.     Equipment was purchased for $17,000 on January 1, 2008. Freight charges amounted to $700 and there was a cost of $2,000 for building a foundation and installing the equipment. It is estimated that the equipment will have a $3,000 salvage value at the end of its 5-year useful life. What is the amount of accumulated depreciation at December 31, 2009, if the straight-line method of depreciation is used?

a.   $6,680

b.   $3,340

c.   $2,860

d.   $5,720
Answer #1

107) B
2007 depreciation would be ($450,000 X 66.67%) = $300,000

2008 depreciation would be ($150,000 X 66.67%) = $100,000

2009 depreciation would be the remaining amount necessary to get the Net Book Value down to the salvage value of $20,000.

In this case, it is $30,000.

108) D
Remaining life = (Depreciable Cost - Book Value) ÷ Annual depreciation = (50,000 - 10,000 - 25,000) ÷ 5,000 = 3 years.

109) B. $11,760.

The amount to be depreciated over the 5-year period is the cost of the asset plus the freight charges plus the construction of the foundation minus the estimated salvage value = $60,000 + $2,800 + $8,000 - $12,000 = $58,800. $58,800/5 years = $11,760 annual depreciation using the straight-line method.

110) A.
Accumulated depreciation= Cost of assest plus freight plus installation expenses minus salvage value/5 years multiplied with number of completed years

$17000+$700+$2000=$19700-$3000=$16700=$16700/5=$3340*2=$6680